OpenStudy (anonymous):

A segment with endpoints A (2, 1) and C (4, 7) is partitioned by a point B such that AB and BC form a 3:2 ratio. Find B. (3.2, 1.6) (3.2, 4.6) (3.8, 3.4) (4.4, 5.8)

OpenStudy (anonymous):

@smiley2015 Can you help with this one?

OpenStudy (anonymous):

@smiley2015

OpenStudy (anonymous):

use segment formula

OpenStudy (anonymous):

What is the segment formula?

OpenStudy (anonymous):

(3*4+2*2)/(3+2) , (3*7+2*1)/(3+2) = 16/5 , 23/5 = (3.2, 4.6)

OpenStudy (anonymous):

if x,y divide the join of the points(x1,y1) and (x2,y2) in the ratio of m:n internally then x=(m x2+n x1)/(m+n) and y =(m y2+n y1)/(m+n)

OpenStudy (anonymous):

Ohh! Thank you!

OpenStudy (anonymous):

yw